How much area under the normal curve lies between the mean and a z-score of 1.30?

0.4032%
40.32%
50%
9.68%

Flag this Question

Question 224 pts

The standardized normal curve always has a mean of _____ and a standard deviation of _____.

50,2
0,1
68,1
1,0.5
Answer #1

Solution :

Using standard normal table,

P( 0 < Z < 1.30) = P(Z < 1.30) - P(Z < 0) = 0.9032 - 0.5 = 0.4032

Area = 40.32%

The standardized normal curve always has a mean of 0 and a standard deviation of 1 .

Answer = (0 , 1)